OpenStudy (anonymous):

dy/dx = (y+8x)/(x+2y) I know the way to the solution is long therefore if you can please tell me what method are you using for the solution and what is the final answer? tnx :)

OpenStudy (anonymous):

btw I used F(y/x)=v(x)

OpenStudy (anonymous):

take x common from Nmr. n Dnr. and put y=vx and then replace dy/dx by v+x(dv/dx) and solve accordingly

OpenStudy (anonymous):

I've done that, that's where things got ugly

OpenStudy (anonymous):

I replaced the dv/dx with dx/dv and isolated the dx/dv on one side minus all the X's and got a Homogeneous equation for x(v)

OpenStudy (anonymous):

btw sry for my lingo, my main language is not English

OpenStudy (anonymous):

I got a solution for x(v) \[x(v)=ce ^{(4-v^2)^(3/8)}\]

OpenStudy (anonymous):

u must have got this\[x(dv/dx)=2(4-v^2)/1+2v\]

OpenStudy (anonymous):

the main problem is that I'm not sure about my final answer

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

that's where I changed the equation from v(x) based to x(v) based. is that what I should have done or is there another way? because when I'm doing this I'm getting a Homogeneous equation which is easy to solve

OpenStudy (anonymous):

after solving u'll get 2+v/(2-v)(4-v^2)^2=x^8c

OpenStudy (anonymous):

ok tnx and if I want to get the : y(x) I need to isolate the v annd put it in the first equation - y(x)=x*v(x) ?

OpenStudy (anonymous):

oh yes to get final answer put v=y/x. Are u in high school?

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

I'm taking a course in Linear differential equation this is from the homework of the second week

OpenStudy (anonymous):

solve this one dy/dx= x+3y+2/3x+9y+6

OpenStudy (anonymous):

sorry this one dy/dx= x+3y+2/3x+9y+5

OpenStudy (anonymous):

dy/dx=1/3 y(x)=1/3*x

OpenStudy (anonymous):

ohh that will take me a few minutes

OpenStudy (anonymous):

-(3x+9y+5)dy+(x+3y+2)dx=0 from here we need to make it into an exact equation

OpenStudy (anonymous):

hey no need for that it can be solve easily

OpenStudy (anonymous):

sry this is an exact equation

OpenStudy (anonymous):

any way Shiv tnx for the help I need to go to another class :)